OpenStudy (anonymous):

Consider an RC circuit with a capacitance of 150 μF and a resistance of 1 MΩ connected to a 200-V power source for a long time, disconnected, and shorted. How long should one wait before the voltage across the capacitor drops to below 50 V? a)208 s b)314 s c)52 s d)104 s e)176 s

OpenStudy (anonymous):

a)208s

OpenStudy (mos1635):

Vc=Vo*e^(-t/RC) set Vc=50 Vo=200 R=10^6 C=150*10^(-6) and solve.......